OpenStudy (hannaha):

Hey, I need some help with this one: if a triangle has 1 3/4, 1 1/4 and 2 than what is the perimeter? Thank you!

hero (hero):

The perimeter of any closed figure is the sum of its side lengths.

hero (hero):

Thus 1 + 3/4 + 1 + 1/4 + 2 = (1 + 1 + 2) + (3/4 + 1/4) = 1 + 1 + 2 + 1